OpenStudy (anonymous):

Find the value of sec theta = 1.263. Give answer in degrees.

OpenStudy (anonymous):

\[\sec \theta = \frac{1}{\cos \theta}\]\[\frac{1}{\cos \theta} = 1.263\] cross multiply and you'll get : \[(\cos \theta)(1.263) = 1\] \[\cos \theta = \frac{1}{1.263}\] now to find the angle in degrees : \[\theta = \cos^{-1} (\frac{1}{1.263}) = 37.6 ^{o}\] crrect me if I'm wrong ^_^